An aqueous solution containing 15.0 g of an unknown molecular (nonelectrolyte) compound in 100.5 g of water has a freezing point of -1.9 ∘C. Calculate the molar mass of the unknown compound. I am stumped because I am not given Kf??
Answer #1

Kf = 1.860C/m

Tf = 0-(-1.9)   = 1.90C

i = 1 for non electrolyte

Tf   = i*Kf*m

1.9    = 1*1.86*m

m   = 1.9/1.86   = 1.0215m

molality   = W*1000/G.M.M.Weight of solvent in g

1.0215   = 15*1000/G.M.M*100.5

G.M.M   = 15*1000/(1.0215*100.5)   = 146g/mole

molar mass of unknown compound = 146g/mole >>>>answer